12.+ SerCP1GE 1P066. fotbai rorn ta od than runssideways acros the feld for 12.5 yards. At this point, he kicks a pass straight upfield to a striker. How far (in yards) is the football from its original location? HINT yd
Answer #1

here,

the the net displacement in forward direction from the original position , y = (54 - 12.5) yd

y = 41.5 yd

the distance travelled across the feild , x = 12 yd

the final distance from the original position , r = sqrt(x^2 + y^2)

r = sqrt(41.5^2 + 12^2) yd

r = 43.2 yd